7. Two 1.80 kg bodies, A and B, collide. The velocities before the collision are and . Immediately after the collision block A has . a. Wat is the velocity of block B immediately after the collision? b. What is the change in the total kinetic energy because of the collision? ((please make the solution readable ))
Answer #1

The speeds are missing, but I'll answer assuming random values (for initial speeds and A speed after the collision):

The conservation of momentum is given by:

m_{1}v_{1} + m_{2}v_{2} = m_{1}v'_{1} + m_{2}v'_{2}

(1.80kg)(2m/s + 4m/s) = (1.80kg)(1.5m/s + v'_{2})

\frac{(1.80kg)(2m/s + 4m/s)}{(1.80kg)} - 1.5m/s = v'_{2}

ANSWER v'_{2} = 4.5m/s

b.-

The change in kinetic energy is:

\Delta KE = \frac{1}{2}m_{1}(v'_{1})^{2} + \frac{1}{2}m_{1}(v'_{2})^{2} - \frac{1}{2}m_{1}(v_{1})^{2} - \frac{1}{2}m_{2}(v_{2})^{2}

\Delta KE = \frac{1}{2}(1.80kg)\left ((1.5m/s)^{2} + (4.5m/s)^{2} - (2m/s)^{2} - (4m/s)^{2} \right )

ANSWER \Delta KE = 2.25J

I hope it helps!!